Write the integer 125 in the manner that every positive integer can be representable as a...

  1. Write the integer 125 in the manner that every positive integer can be representable as a sum of Fibonacci numbers, none taken more than once. Example: 5 = F3 + F4
Answer #1

# Fibonacci numbers. fn=foritfn-2. hi fa fa fy is to f fe fy fio fin 1 2 3 5 8 13 4 34 55 89-.. 195= 2 + 34+89 1185= fg + f g